OpenStudy (anonymous):

A red ball and 6 white balls are in a box. If two balls are drawn, without replacement, what is the probability of getting a red ball on the first draw and a white ball on the second? a. 1/6 b. 1/7 c. 6/7 d. 7/13

OpenStudy (anonymous):

You have a 1/7 chance to draw a red ball first. Since you aren't replacing the ball or putting it back into the box, all you have left are white balls in the box. So you have a 100% or 1.0 chance to get the white ball. So: \[1/7 \times 1 = 1/7\]

OpenStudy (anonymous):

what about the white ball on the second draw?

OpenStudy (anonymous):

You have a 100% chance of getting the white ball. So its 1/7 multiplied by 1. The 1 is the 100%.

OpenStudy (anonymous):

1/7 the probability of red coming for the first time is 1/7 and then there are only white ones whose probability become 1

OpenStudy (anonymous):

Notation wise, it can be written as: \[(1/7) \times (6/6)\]

OpenStudy (anonymous):

Which is equal to \[1/7\]

OpenStudy (anonymous):

oooh i see. thank you guys!